OpenStudy (anonymous):

A = P + I = P + Prt = P (1+rt) ?Why are there multiple versions for finding A?

Parth (parthkohli):

Because there are many ways to solve a problem. But all of them mean the same thing.

Parth (parthkohli):

P + I = P + PRT because I = PRT.

Parth (parthkohli):

Just apply yourself and see what you are comfortable with.

OpenStudy (anonymous):

well if they all mean the same i get different conclusions with them I'll show you what I mean.

OpenStudy (anonymous):

For instance, A person borrows $ 1000 for a period of 6 months. What simple interest rate is being charged if the amount A that must be repaid after 6 months is $ 1045.

OpenStudy (anonymous):

The formula in textbook used A = P + Prt so that is 1045 = 1000 +1000r(1/2) ... that leads to 45 = 500r and then r becomes...09. However if i used A = P ( 1+rt) the awnser i got was different.

OpenStudy (anonymous):

i got.. r = 1045/1500..which means r = .69...

OpenStudy (anonymous):

D: i don't understand...

Parth (parthkohli):

\[A = P(1 + RT)\]\[\Rightarrow 1045 = 1000(1 + 0.5R)\]\[\Rightarrow 1045= 1000 + 500R\]\[\Rightarrow500R = 45\]

Parth (parthkohli):

Also, \(P+PRT = P(1 + RT)\) because we have factored the \(P\) out.

OpenStudy (anonymous):

wait.. shouldn't (1+0.5) 1.5 ... 1.5 (1000) is 1500? how did you get 500?

OpenStudy (anonymous):

i thought it was 1045 = 1500r =T

Parth (parthkohli):

That is \(0.5R\), not \(0.5\). We know that \(T =0.5\).\[P(1 + RT) = P(1 + 0.5R) = 1000(1 + 0.5R) = 1000 + 500R\]

OpenStudy (anonymous):

sorry.. i still don't understand.. T_T... for some reason i can't get 500r T_T

OpenStudy (anonymous):

i feel so stupid..T_T

Parth (parthkohli):

\[1000(1 + 0.5R)\]Do you get it till here?

OpenStudy (the_fizicx99):